Question:- If | | = | | = | | then angle between A and B will be: (a) 90 0 (b) 120 0 (c) 0 0 (d) 60 0 Plzz tell me with procedure hw to solve this

 

Question:-

If |\vec{A}+\vec{B}| = |\vec{A}| = |\vec{B}| then angle between A and B will be:

(a) 90      (b) 1200        (c) 00    (d) 600                                   Plzz tell me with procedure hw to solve this

Grade:Select Grade

3 Answers

Akshat Jain
9 years ago
It is 120. 
Use the formula of vector sum by using |\vec{A}| = |\vec{B}|
and angle between them 120
 
 
 
Mohammad Kavish
8 years ago
Akshat I dnt get it
noogler
8 years ago
given Ia+bI=IaI=IbI
on squaring it
Ia+bI2=IaI2=IbI2
IaI2+IbI2+2IaIIbIcoso=IaI2
2IaI2+2IaI2coso=IaI2 ….....................................since IaI=IbI
2(1+coso)=1
coso=-1/2................................o=1/2
o is angle btw a and b